Form the quadratic equation for which the sum of the roots is 5 and the sum of the squares of the roots is 53

1
Expert's answer
2022-02-02T08:23:28-0500

Let the two roots be "\alpha" and "\beta".

"\alpha +\beta=5\\
\alpha ^2+\beta^2=53"

But,

"(\alpha+\beta)^2=\alpha^2+\beta^2+2\alpha \beta\\
2\alpha \beta=(\alpha+\beta)^2-\alpha^2-\beta^2\\
2\alpha \beta=5^2-53=-28\\
\alpha \beta=-14"

Quadratic equation with two roots "\alpha" and "\beta" is given as:

"x^2-(\alpha+\beta)x+\alpha \beta=0\\
\text{Substitute in the unknowns, we have the quadratic equation:}\\
x^2-5x-14=0"